Pix Company has the following production data for March: no beginning work in process, units started and completed 29,500, and ending work in process 5,100 units that are 100% complete for materials and 40% complete for conversion costs. Pix uses the FIFO method to compute equivalent units. If unit materials cost is $5 and unit conversion cost is $14, determine the costs to be assigned to the units transferred out and the units in ending work in process. The total costs to be assigned are $614,560.

Transferred out

$

Ending work in process

$
Answer #1

B с D E $ $ 560,500 54,060 A 1 Transferred out 2 Ending work in process 3 4 Working note: 5 6 Quantities 7 Units accounted fo

B с D E =+E25 =+E28 Physical units Materials Conversion costs 29500 5100 =+B8*100% =+B8*100% =+B9*100% =+B9*40% =+C8+C9 =+D8+